OpenStudy (theedwardsfamily):

Which answer best describes the solution to the equation 3m – 9 = 9 + 3m? A. one solution B. no solution C. an infinite number of solutions D. m = 27

OpenStudy (anonymous):

B

OpenStudy (theedwardsfamily):

plz explain why...

OpenStudy (anonymous):

Alright so what you do is combine "like terms" 3m and 3m will be next to each other and the 9's will be next to each other: \[3m - 3m = 9 + 9\] \[0 = 18\] So yes, B is correct.

OpenStudy (theedwardsfamily):

ok tnx

OpenStudy (anonymous):

Remember that you can NOT give out answers @tokidoki007. Please read this ---> http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

And np @TheEdwardsFamily :)

OpenStudy (theedwardsfamily):

ik ,lag